# Triad: A Framework Leveraging a Multi-Role LLM-based Agent to Solve Knowledge Base Question Answering

### v4

@@ Abstract
Chang Zong Zhejiang University zongchang@zju.edu.cn                        Yuchen Yan Zhejiang University yanyuchen@zju.edu.cn                        Weiming Lu Zhejiang University luwm@zju.edu.cn                        Jian Shao Zhejiang University jshao@zju.edu.cn
Eliot Huang Leibowitz AI EliotHuang@leibowitz.org.cn                        Heng Chang Tsinghua University changh17@tsinghua.org.cn                        Yueting Zhuang Zhejiang University yzhuang@zju.edu.cn
@@ 1 Introduction
In this study, we introduce Triad, a unified framework that leverages an LLM-based agent with three roles to address KBQA tasks. Specifically, we implement the agent consisting of an LLM as the core, supplemented by various task-specific modules such as memory and executing functions. The agent is assigned three distinct roles: a generalist (G-Agent) adept at mastering numerous small tasks by the given examples, a decision maker (D-Agent) proficient at identifying options and selecting candidates, and an advisor (A-Agent) skilled at providing answers using internal and external knowledge. The cooperation of these agent roles composes a KBQA process containing four phases: question parsing, URI linking, query construction, and answer generation. We evaluate our framework on three benchmark datasets in various difficulties. The results show that our framework outperforms state-of-the-art systems, demonstrated by 11.8% and 20.7% F1 scores on the LC-QuAD and YAGO-QA benchmarks, respectively.
The contributions of this study can be summarized as follows: (1) We propose Triad, the first framework that leverages an LLM-based agent to solve KBQA tasks in all its four phases, without specialized training models. (2) We implement an LLM-based agent with various task-specific modules that can act as three roles, including a generalist, a decision maker, and an advisor, to collaboratively solve KBQA via focusing on subtasks. (3) We evaluate the performance of Triad. The results show a competitive ability compared to both state-of-the-art KBQA systems and pure LLM methods [^1] Code and datasets used to achieve the performance will be made public if the paper is accepted..
@@ 3.1 G-Agent as a Generalized Solver
$Pmt_{qt}=$ $\left[Ins_{qt},Shot_{qt},CoT_{qt}\right],\theta_{qt}=\left[Q,f(S_{tri})\right]$
@@ 4.3 Study on Capabilities of Agent Roles
We assess the efficacy of G-Agent with various other language models as the core. The framework without G-task uses the text-davinci-002, which is not as powerful as GPT-3.5 and GPT-4 in solving many tasks, and the one without G-chat uses text-davinci-003 to eliminate the chat and alignment abilities. We test the ability of D-Agent without D-uri and D-query by replacing the URI selection and query selection with URI matching and query generation, respectively. We evaluate the contribution of A-Agent eliminating A-llm and A-fact by responding to questions without using LLM’s assistance or use an LLM to answer Boolean questions for auxiliary rather than single-fact questions. The F1 results of the role ablation experiments are shown in Table 3. The results indicate that every component pertaining to each role contributes to the overall performance. More specifically, a G-Agent that employs a less powerful LLM as its core can drastically undermine performance. D-Agent assumes a more pivotal role during the linking phase compared to the query construction phase. A-Agent, on the other hand, proves to be an efficient solution for managing situations where SPARQL results are absent.
@@ 6 Conclusion
The limitation of our research lies in a further research and discussion on a more diverse range of multi-agent collaboration manners for KBQA tasks. Moreover, this framework worth evaluating on domain specific QA datasets.
@@ Appendix A Response Time Analysis
Appendix B Prompts Provided to LLMs of G-Agent for Solving Various Subtasks in KBQA
@@
Appendix C Prompts Provided to LLMs of D-Agent for Solving Selection Subtasks in KBQA
@@
Appendix D Prompts Provided to LLMs of A-Agent for Solving Answering Subtask in KBQA

### v6

@@ Abstract
[…]
@@ 1 Introduction
In this study, we introduce Triad, a unified framework that leverages an LLM-based agent with three roles to address KBQA tasks. Specifically, we implement the agent consisting of an LLM as the core, supplemented by various task-specific modules such as memory and executing functions. The agent is assigned three distinct roles: a generalist (G-Agent) adept at mastering numerous small tasks by the given examples, a decision maker (D-Agent) proficient at identifying options and selecting candidates, and an advisor (A-Agent) skilled at providing answers using internal and external knowledge. The cooperation of these agent roles composes a KBQA process containing four phases: question parsing, URI linking, query construction, and answer generation. We evaluate our framework on three benchmark datasets in various difficulties. The results show that our framework outperforms state-of-the-art systems, demonstrated by 11.8% and 20.7% F1 scores on the LC-QuAD and YAGO-QA benchmarks, respectively [^1] Code and data are available at https://github.com/ZJU-DCDLab/Triad..
The contributions of this study can be summarized as follows:
  • We propose Triad, the first framework that leverages an LLM-based agent to solve KBQA tasks in all its four phases, without specialized training models.
  • We implement an LLM-based agent with various task-specific modules that can act as three roles, including a generalist, a decision maker, and an advisor, to collaboratively solve KBQA via focusing on subtasks.
  • We evaluate the performance of Triad. The results show a competitive ability compared to both state-of-the-art KBQA systems and pure LLM methods.
@@ 3.1 G-Agent as a Generalized Solver
$Pmt_{qt}=$ $\left[Ins_{qt},Shot_{qt},CoT_{qt}\right],$
$\theta_{qt}=$ $\left[Q,f(S_{tri})\right]$
@@ 4.3 Study on Capabilities of Agent Roles
We assess the efficacy of G-Agent with various other language models as the core. The framework without G-task uses the text-davinci-002, which is not as powerful as GPT-3.5 and GPT-4 in solving many tasks, and the one without G-chat uses text-davinci-003 to eliminate the chat and alignment abilities. We test the ability of D-Agent without D-uri and D-query by replacing the URI selection and query selection with URI matching and query generation, respectively. We evaluate the contribution of A-Agent eliminating A-llm and A-fact by responding to questions without using LLM’s assistance or use an LLM to answer Boolean questions for auxiliary rather than single-fact questions. The F1 results of the role ablation experiments on two representative datasets are shown in Table 3. The results indicate that every component pertaining to each role contributes to the overall performance. More specifically, a G-Agent that employs a less powerful LLM as its core can drastically undermine performance. D-Agent assumes a more pivotal role during the linking phase compared to the query construction phase. A-Agent, on the other hand, proves to be an efficient solution for managing situations where SPARQL results are absent.
@@ 4.6 Study on Complex Cases
4.7 Cost Comparison and Analysis
According to our evaluation on the three datasets, the average cost of running a single case is 0.007 USD on average using Triad-GPT3.5 and 0.05 USD on average using Triad-GPT4. Specifically, most API calls occur in the phases of URL linking and comprehensive answering. Meanwhile, traditional KBQA baselines require a lot of training data and local training resources to achieve the SOTA performance, whereas Triad follows a zero- or few-shot manner to save computational cost locally. Furthermore, as shown in Section 4.4, in practice, adjusting the hyperparameters can make the cost as low as possible while preserving overall performance. As the cost of LLM services decreases, the value of Triad will increase accordingly.
@@ 6 Conclusion
The limitation of our research lies in following aspects: (1) In terms of data, a broader range of QA datasets needs to be evaluated, encompassing datasets from different domains, languages, and difficulty levels. (2) In terms of model, more LLMs need to be evaluated, including open-source and commercial models from different organizations and on various scales. (3) In terms of framework, more types of agent collaboration methods can be explored to solve KBQA problems.
@@
Acknowledgements
This work is supported by the "Pioneer" and "Leading Goose" R&D Program of Zhejiang (Grant No. 2023C01152).
@@ Appendix A Response Time Analysis
Appendix B Role Performance on YAGO-QA
We choose LC-QuAD 1.0 and QALD-9 as our two representative datasets in Section 4.3, as the questions among them vary in difficulty, and the tasks in these two datasets are relatively more challenging than YAGO-QA. We provide the performance of agent roles on YAGO-QA in Table 6, which shows a consistent result with other datasets in Table 3.
G-task G-chat YAGO-QA
0.427
0.553
0.677
D-uri D-query YAGO-QA
0.346
0.534
0.677
A-llm A-fact YAGO-QA
0.626
0.647
0.677
Table 6: Performance of roles of LLM-based agent by eliminating an element or downgrading the capability.
Appendix C Prompts Provided to LLMs of G-Agent for Solving Various Subtasks in KBQA
@@
Appendix D Prompts Provided to LLMs of D-Agent for Solving Selection Subtasks in KBQA
@@
Appendix E Prompts Provided to LLMs of A-Agent for Solving Answering Subtask in KBQA